The gain of an amplifier is found to be G=20log(10Vout):
The tasks are to find equations for :
A. dG/dVout
B. d^2G/dVout^2
1
Expert's answer
2020-03-26T14:42:38-0400

"\frac{\partial G}{\partial Vout}=\frac{20}{ln10*Vout}, 
\frac{\partial^2 G}{\partial^2 Vout} =\frac{-20}{ln10}*\frac{1}{Vout^2}"